OpenStudy (anonymous):

the cost in dollars n for manufacturing t refrigerator makers is given by the following equation n=8t+65. how many waffle makers can be manufactured for a total of &2575.

OpenStudy (nikato):

i think you made some typos so the total cost is 2575. plug that into n in the equation so 2575=8t+65 now solve for t

OpenStudy (anonymous):

oh yes...how many refrigerators can be manufactured for a total of 2575

OpenStudy (nikato):

ok, so can you solve for t?

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

2575=8t+65 =2575/8=8t/8+65/8

OpenStudy (nikato):

nice try. but using PEMDAS, u would go backward, so your first step would be subtracting 65 on both sides 2575=8t+65 -65 -65 ------------- 2510=8t ---- -- 8 8 t=?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

313.75

OpenStudy (nikato):

yes and you would round down to 313 because you cant have .75 of a refrigerator and you cant round up, because you wont have enough money to make that remaining .25 of a refrigerator

OpenStudy (anonymous):

thanks

OpenStudy (nikato):

no problem